# Does VLM Classification Benefit from LLM Description Semantics?

### v2

@@ Abstract
Codehttps://github.com/CompVis/DisCLIP
Extended Versionhttps://arxiv.org/abs/2412.11917
@@ 1 Introduction
Vision-Language Models (VLM) (Radford et al. 2021; Jia et al. 2021) tackle this issue by training neural networks to link images and their textual descriptions within a shared embedding space. This enhances the correlation between visual and textual details. VLMs can be applied to zero-shot image classification by passing an image through the VLM’s image encoder and prompting the text encoder with hand-crafted inputs like “a photo of a [classname]” (Radford et al. 2021). Recent work(Menon and Vondrick 2022) extends this approach by incorporating additional descriptions generated by Large Language Models (LLMs) for each class name. LLMs like GPT-3 (Brown et al. 2020; Ouyang et al. 2022) or Llama (Touvron et al. 2023), trained on extensive text corpora, are intended to provide richer semantics, enhancing VLM classification.
@@ 2 Related Works
Vision-language models such as (Radford et al. 2021) can be used for classification. Notable training-free approaches that build on top of this include DCLIP (Menon and Vondrick 2022) and CuPL (Pratt et al. 2023), where class name texts are augmented with knowledge contained in LLMs to leverage seemingly discriminating characteristics to achieve performance boosts. As (Roth et al. 2023) demonstrated, similar effects could be obtained by augmenting the class name with text noise and high-level concepts, raising concerns that many performance gains from DCLIP (Menon and Vondrick 2022) were not due to additional semantics but rather to introduced noise. FuDD (Esfandiarpoor and Bach 2024) introduced contrastive zero-shot prompting to obtain a more diverse set of text prompts. The disadvantage of these approaches is that they rely on ensembling the description extended class name multiple times to achieve significant gains, making it difficult to separate additional semantics from random augmentations of the class name.
Notable approaches that train in the CLIP embedding space include (Yan et al. 2023), where nearest neighbors from a pool of text embeddings replace linear weights of a learned dictionary, and LaBo (Yang et al. 2023), which trained a linear classifier on a wide and global bottleneck of language activations selected for diversity and coverage. (Zhou et al. 2022) performed non-explainable tuning of text prompt embeddings to optimize classification, while (Zang et al. 2024) trained the last layer of image and text encoders over a concept bottleneck to discover explainable concepts. (Feng, Bair, and Kolter 2023) trained a sparse logistic regression over a matrix of image-language activations, with the training signal also used to train the image encoder.
In contrast to the methods outlined above, our approach delivers humanly understandable (thus explainable), semantically meaningful, disjoint, and distinguishing language descriptions in text space through a training-free method, which boosts VLMs classification accuracy while providing higher explainability. Reasons for this better explainability can be found in Section 3.2.
@@
Test Time Augmentation TTA introduces an ensemble of predictions from several transformed or distorted versions of a given test input to obtain a “smoothed” prediction. For example, one could average the predictions from various modified versions of a given string, ensuring that the final prediction is robust to any single unfavorable version (Roth et al. 2023; Menon and Vondrick 2022; Esfandiarpoor and Bach 2024).
Some previous methods (Esfandiarpoor and Bach 2024) used up to hundreds of thousands of descriptions per class, achieving significant improvements in classification accuracy with VLM. However, it is challenging to determine if the performance gains result from the vast ensemble or true information, hence hindering explainability.
@@ 3.1 Background
In the course of this work and w.r.t. to textual descriptions, we call a set $\mathcal{D}(c)$ a description assignment. Furthermore, an LLM ”assigns” descriptions by generating them when prompted for a particular class $c$, yielding the assignments $D(c)$. The elements of $\mathcal{D}(c)$ can be pure text augmentations, e.g. “an image of [cls], “a photo of [cls], or can contain LLM-generated text descriptions and high-level-concepts, e.g. “an image of [cls], a type of [LLM-generated category], with [LLM-generated descriptions]. Most of the approaches (Menon and Vondrick 2022; Pratt et al. 2023; Esfandiarpoor and Bach 2024; Roth et al. 2023) that use ensembling as in Equation 3 with LLM-generated contents always include the class name token [cls] for $\forall d\in\mathcal{D}(c)$, which we denoted as classname-included descriptions.
@@ 3.2 Our Approach
In the conventional setup (Menon and Vondrick 2022; Pratt et al. 2023; Esfandiarpoor and Bach 2024), performance gains may result from the noise augmentation of the class name text [cls] embedding through its various combinations with [LLM-generated category], [LLM-generated descriptions], and even random strings. While random strings should not contribute extra semantics and are likely embedded far away from [cls], this can sometimes apply to LLM-assigned text due to the vocabulary discrepancy between VLMs and LLMs. Another cause may also be that the images do not exhibit the described property. Despite this, such combinations can still perform well, although the assigned descriptions are not semantically correct.
@@
The selected descriptions can then be used as described in Section 3.2 or Section 3.1 for inference. In both cases, classification happens via an ensembling (classname-containing and classname-free) of image-language similarities, as introduced in (Radford et al. 2021). Applying $\operatorname*{arg\,max}$ over the ensembled, description-enriched image-language similarities of the candidate set $\mathcal{A}(x_{i})$ yields the final classification decision of the image.
@@
    The number of resulting textual descriptions for a single class is kept within a reasonable limit, similar to the approach in the seminal work (Menon and Vondrick 2022). This helps minimize the potential for noise augmentation, unlike methods that generate hundreds and thousands of descriptions (Esfandiarpoor and Bach 2024).
@@
    We do not use continuous weights over resulting textual descriptions, as done in  (Yang et al. 2023; Yan et al. 2023; Zang et al. 2024). Long vectors of continuous weights can be less interpretable compared to clear, discrete indicators of whether a concept is present. Hence, our method offers improved clarity and explainability.
@@ 4.1 Implementation Details and Datasets
Implementation details. We use CLIP (Radford et al. 2021) as the base Vision-Language Model (VLM) for our approach. Unless stated otherwise, the backbone for CLIP is the ViT-B/32 (Vaswani et al. 2017; Dosovitskiy et al. 2020). We randomly sample a subset from each dataset’s standard training split to obtain the lookup similarity table S (details see Section A.9). Our empirical tests confirmed that this sampling process does not significantly impact performance. The Large-Language Model (LLM) generated descriptions are sourced directly from DCLIP (Menon and Vondrick 2022) or generated using the contrastive prompting method with gpt-3.5-turbo-1106 and Llama-3-70b-chat-hf via APIs.
@@
We evaluated our methods on the following standard datasets using the standard protocol (classification accuracy) based on previous works (Menon and Vondrick 2022; Roth et al. 2023): ImageNet (Deng et al. 2009), ImageNetV2 (Recht et al. 2019), CUB200-2011 (Wah et al. 2011) (fine-grained bird classification), EuroSAT (Helber et al. 2017) (satellite image recognition), Places365 (Zhou et al. 2017), DTD (Textures, (Cimpoi et al. 2014)), and Flowers102 (Nilsback and Zisserman 2008).
@@
These descriptions can be obtained in the following ways: 1) directly from the published descriptions of other works, such as DCLIP  (Menon and Vondrick 2022) or FUDD  (Esfandiarpoor and Bach 2024); 2) generated based on the provided procedures and code bases of other works, if descriptions are not available; 3) or created through contrastive prompting, which aims to extract meaningful descriptions by contrasting hard negative samples within a neighborhood. The motivation is similar to FuDD  (Esfandiarpoor and Bach 2024), but we use significantly fewer descriptions per class. As this is only an alternative for constructing a description pool and orthogonal to our proposed method, we provide more details in Section A.7 on the construction of the contrastive pool.
@@ 5 Conclusion
This study found that VLM Classification performance does indeed benefit from LLM description semantics - if the descriptions are correctly selected. For that, we present a new, training-free approach that assigns semantically meaningful descriptions based on feedback from the VLM embedding space. We show that descriptions assigned in this way have discriminative power on their own, i.e. when evaluated in our proposed setup without classname ensembling. These description assignments improve the performance in image classification tasks - with and without classname ensembling. Furthermore, this evaluation setup helps to differentiate performance gains from genuine semantics from gains from ensemble effects. Hopefully, our findings inspire future work on VLMs and model explainability.
@@ A.2 Different choices of size $k$ and $n$
Table 4 ablates the number $k$ of preliminary labels per test image and $n$, the number of images per class available in the selection set for S. Results were obtained in the classname-included evaluation setup on the CUB200 Dataset with $5$ selected descriptions from the DClip description pool. Increasing the number of selection samples in $S$ leads to increasing performances. For the CUB dataset, the optimal $k$ value is empirically found to be $k=3$ as it has a right trade-off between a diverse set of candidates that is still small enough to find distinctive descriptions. This shows that even in a few-shot regime with only $5$ selected descriptions per candidate class baselines such as  (Esfandiarpoor and Bach 2024) and  (Menon and Vondrick 2022) can be exceeded (cf.  Table 3 in the main text).
@@ A.6 Evaluating Beyond Ensembling
To construct the description pool $\mathcal{P}$, one source is to utilize published descriptions from other works, such as DCLIP (Menon and Vondrick 2022). However, since DCLIP does not provide descriptions for Flowers102, we generated descriptions for this dataset using GPT-3.5 with prompts from their code base. In these scenarios, the LLM ”assigns” descriptions to a class by generating them specifically for that class. Therefore, both LLM assignments and our assignments, as described in  algorithm 1, are included in the DCLIP description pool. In rare instances where the LLM did not return a description following the DCLIP approach, a neutral description, such as ”a kind of food” for the food dataset, was used. This step was necessary to avoid distorting the results in this setup otherwise.
@@ A.8 Comparison to prior works in the classname-free setting
  • Works like LaBo (Yang et al. 2023) use large Bottleneck sizes (from 500 for CUB up to 50,000 for ImageNet. They are mostly larger than the original CLIP embedding dimensionality, cf. Table 16 in the appendix of (Yang et al. 2023)). Therefore, the CLIP image embedding gets transformed to an overcomplete basis and overlaps 100% between classes. This requires that several thousands of continuous linear classifier weights be interpreted for every class, which is quite cumbersome.
@@
(Vogel et al. 2022) evaluated VLM classification by relying only on language descriptions but did not analyze the interplay between separate descriptions and a simultaneously introduced classname prompt $d^{c+}$, which is crucial to understand the resulting accuracy in VLM ensembling scenarios. In addition to that, every image description was appended to the same text prompt. The resulting long prompts degraded performance.

### v3

@@ 1 Introduction
Vision-Language Models (Radford et al. 2021; Jia et al. 2021) tackle this issue by training neural networks to link images and their textual descriptions within a shared embedding space. This enhances the correlation between visual and textual details. VLMs can be applied to zero-shot image classification by passing an image through the VLM’s image encoder and prompting the text encoder with hand-crafted inputs like “a photo of a [classname]” (Radford et al. 2021). Recent works (Menon and Vondrick 2023; Chiquier, Mall, and Vondrick 2024) extends this approach by incorporating additional descriptions generated by Large Language Models (LLMs) for each class name. LLMs like GPT-3 (Brown et al. 2020; Ouyang et al. 2022) or Llama (Touvron et al. 2023), trained on extensive text corpora, are intended to provide richer semantics, enhancing VLM classification.
@@ 2 Related Works
Vision-language models such as (Radford et al. 2021) can be used for classification. Notable training-free approaches that build on top of this include DCLIP (Menon and Vondrick 2023) and CuPL (Pratt et al. 2023), where class name texts are augmented with knowledge contained in LLMs to leverage seemingly discriminating characteristics to achieve performance boosts. As Roth et al. (2023) demonstrated, similar effects could be obtained by augmenting the class name with text noise and high-level concepts, raising concerns that many performance gains from DCLIP (Menon and Vondrick 2023) were not due to additional semantics but rather to introduced noise. FuDD (Esfandiarpoor and Bach 2024) introduced contrastive zero-shot prompting to obtain a more diverse set of text prompts. The disadvantage of these approaches is that they rely on ensembling the description extended class name multiple times to achieve significant gains, making it difficult to separate additional semantics from random augmentations of the class name.
Notable approaches that train in the CLIP embedding space include Yan et al. (2023), where nearest neighbors from a pool of text embeddings replace linear weights of a learned dictionary, and LaBo (Yang et al. 2023), which trained a linear classifier on a wide and global bottleneck of language activations selected for diversity and coverage. Zhou et al. (2022) performed non-explainable tuning of text prompt embeddings to optimize classification. In contrast, Zang et al. (2024) trained the last layer of image and text encoders over a concept bottleneck to discover explainable concepts. Feng, Bair, and Kolter (2023) trained a sparse logistic regression over a matrix of image-language activations, with the training signal also used to train the image encoder.
In contrast to the methods outlined above, our approach delivers humanly understandable (thus explainable), semantically meaningful, disjoint, and distinguishing language descriptions in text space through a training-free method. It boosts VLMs classification accuracy while providing higher explainability. We will further discuss the better explainability in Section 3.2.
@@
Test Time Augmentation TTA introduces an ensemble of predictions from several transformed or distorted versions of a given test input to obtain a “smoothed” prediction. For example, one could average the predictions from various modified versions of a given string, ensuring that the final prediction is robust to any single unfavorable version (Roth et al. 2023; Menon and Vondrick 2023). On the other hand, Esfandiarpoor and Bach (2024) used up to hundreds of thousands of descriptions per class, achieving significant improvements in classification accuracy with VLM. However, it is challenging to determine if the performance gains result from the vast ensemble or true information, hence hindering explainability.
@@ 3.1 Background
In the course of this work and w.r.t. to textual descriptions, we call a set $\mathcal{D}(c)$ a description assignment. Furthermore, an LLM ”assigns” descriptions by generating them when prompted for a particular class $c$, yielding the assignments $D(c)$. The elements of $\mathcal{D}(c)$ can be pure text augmentations, e.g. “an image of [cls], “a photo of [cls], or can contain LLM-generated text descriptions and high-level-concepts, e.g. “an image of [cls], a type of [LLM-generated category], with [LLM-generated descriptions]. Most of the approaches (Menon and Vondrick 2023; Pratt et al. 2023; Esfandiarpoor and Bach 2024; Roth et al. 2023) that use ensembling as in Equation 3 with LLM-generated contents always include the class name token [cls] for $\forall d\in\mathcal{D}(c)$, which we denoted as classname-included descriptions.
@@ 3.2 Our Approach
In the conventional setup (Menon and Vondrick 2023; Pratt et al. 2023; Esfandiarpoor and Bach 2024), performance gains may result from the noise augmentation of the class name text [cls] embedding through its various combinations with [LLM-generated category], [LLM-generated descriptions], and even random strings. While random strings should not contribute extra semantics and are likely embedded far away from [cls], this can sometimes apply to LLM-assigned text due to the vocabulary discrepancy between VLMs and LLMs. Another cause may also be that the images do not exhibit the described property. Despite this, such combinations can still perform well, although the assigned descriptions are not semantically correct.
@@
The selected descriptions can then be used as described in Section 3.2 or Section 3.1 for inference. In both cases, classification happens via an ensembling (classname-containing and classname-free) of image-language similarities, as introduced in DCLIP. Applying $\operatorname*{arg\,max}$ over the ensembled, description-enriched image-language similarities of the candidate set $\mathcal{A}(x_{i})$ yields the final classification decision of the image.
@@
    The number of resulting textual descriptions for a single class is kept within a reasonable limit, similar to the approach in the seminal work (Menon and Vondrick 2023). This helps minimize the potential for noise augmentation, unlike methods that generate hundreds and thousands of descriptions (Esfandiarpoor and Bach 2024).
@@
    We do not use continuous weights over resulting textual descriptions, as done in  Yang et al. (2023); Yan et al. (2023); Zang et al. (2024). Long vectors of continuous weights can be less interpretable compared to clear, discrete indicators of whether a concept is present. Hence, our method offers improved clarity and explainability.
@@ 4.1 Implementation Details and Datasets
Implementation details. We use CLIP (Radford et al. 2021) as the base Vision-Language Model (VLM) for our approach. Unless stated otherwise, the backbone for CLIP is the ViT-B/32 (Vaswani et al. 2017; Dosovitskiy et al. 2021). We randomly sample a subset from each dataset’s standard training split to obtain the lookup similarity table S (details see Section A.9). Our empirical tests confirmed that this sampling process does not significantly impact performance. The Large-Language Model (LLM) generated descriptions are sourced directly from DCLIP (Menon and Vondrick 2023) or generated using the contrastive prompting method with gpt-3.5-turbo-1106 and Llama-3-70b-chat-hf via APIs.
@@
We evaluated our methods on the following standard datasets using the standard protocol (classification accuracy) based on previous works (Menon and Vondrick 2023; Roth et al. 2023): ImageNet (Deng et al. 2009), ImageNetV2 (Recht et al. 2019), CUB200-2011 (Wah et al. 2011) (fine-grained bird classification), EuroSAT (Helber et al. 2017) (satellite image recognition), Places365 (Zhou et al. 2017), DTD (Textures, (Cimpoi et al. 2014)), and Flowers102 (Nilsback and Zisserman 2008).
@@
These descriptions can be obtained in the following ways: 1) directly from the published descriptions of other works, such as DCLIP  (Menon and Vondrick 2023) or FUDD  (Esfandiarpoor and Bach 2024); 2) generated based on the provided procedures and code bases of other works, if descriptions are not available; 3) or created through contrastive prompting, which aims to extract meaningful descriptions by contrasting hard negative samples within a neighborhood. The motivation is similar to FuDD  (Esfandiarpoor and Bach 2024), but we use significantly fewer descriptions per class. As this is only an alternative for constructing a description pool and orthogonal to our proposed method, we provide more details in Section A.7 on the construction of the contrastive pool.
@@ 5 Conclusion
This study demonstrates that VLM Classification performance indeed benefits from LLM description semantics - if the descriptions are correctly selected. To achieve this, we introduce a training-free method that assigns semantically meaningful descriptions based on feedback from the VLM embedding space. Our results indicate that these descriptions possess inherent discriminative power, as evidenced by evaluations conducted without classname ensembling in our proposed setup. Furthermore, incorporating these description assignments enhances performance in image classification tasks, both with and without classname ensembling. Additionally, our evaluation framework effectively distinguishes performance improvements arising from genuine semantic understanding from those resulting from ensemble effects. We hope that our findings will inspire future research on VLMs and contribute to the development of models with enhanced explainability.
@@ A.2 Different choices of size $k$ and $n$
Table 4 ablates the number $k$ of preliminary labels per test image and $n$, the number of images per class available in the selection set for S. Results were obtained in the classname-included evaluation setup on the CUB200 Dataset with $5$ selected descriptions from the DClip description pool. Increasing the number of selection samples in $S$ leads to increasing performances. For the CUB dataset, the optimal $k$ value is empirically found to be $k=3$ as it has a right trade-off between a diverse set of candidates that is still small enough to find distinctive descriptions. This shows that even in a few-shot regime with only $5$ selected descriptions per candidate class baselines such as  Esfandiarpoor and Bach (2024) and  Menon and Vondrick (2023) can be exceeded (cf.  Table 3 in the main text).
@@ A.6 Evaluating Beyond Ensembling
To construct the description pool $\mathcal{P}$, one source is to utilize published descriptions from other works, such as DCLIP (Menon and Vondrick 2023). However, since DCLIP does not provide descriptions for Flowers102, we generated descriptions for this dataset using GPT-3.5 with prompts from their code base. In these scenarios, the LLM ”assigns” descriptions to a class by generating them specifically for that class. Therefore, both LLM assignments and our assignments, as described in  Algorithm 1, are included in the DCLIP description pool. In rare instances where the LLM did not return a description following the DCLIP approach, a neutral description, such as ”a kind of food” for the food dataset, was used. This step was necessary to avoid distorting the results in this setup otherwise.
@@ A.8 Comparison to prior works in the classname-free setting
  • Works like LaBo (Yang et al. 2023) use large Bottleneck sizes (from 500 for CUB up to 50,000 for ImageNet. They are mostly larger than the original CLIP embedding dimensionality, cf. Table 16 in the appendix of Yang et al. (2023). Therefore, the CLIP image embedding gets transformed to an overcomplete basis and overlaps 100% between classes. This requires that several thousands of continuous linear classifier weights be interpreted for every class, which is quite cumbersome.
@@
Vogel et al. (2022) evaluated VLM classification by relying only on language descriptions but did not analyze the interplay between separate descriptions and a simultaneously introduced classname prompt $d^{c+}$, which is crucial to understand the resulting accuracy in VLM ensembling scenarios. In addition to that, every image description was appended to the same text prompt. The resulting long prompts degraded performance.